# TetSphere Splatting: Representing High-Quality Geometry with Lagrangian Volumetric Meshes

### v2

@@ Appendix 0.F Related work – Text-to-3D content generation
The recent success of text-to-2D image generation models has spurred a growing interest in generating 3D output from text input. In light of the limited availablity of text-annotated 3D datasets, methods have been developed that leverage the pre-trained text-to-image models to reason between 2D renderings of 3D models and text descriptions. Early works [42, 35] adopt the pre-trained CLIP model [70] to supervise the generation by aligning the clip text and image embeddings. More recently, 2D diffusion-based generative models [73, 71] have powered direct supervision in the image/latent space and achieved superior 3D quality [67, 84, 11, 46, 93]. Notably, DreamFusion [67] introduces the Score Distillation Sampling (SDS) for supervising the NeRF [58] optimization using diffusion priors as a score function (i.e., by minimizing the added noise and the predicted noise under the text condition). Follow-up works have since been proposed to improve score sampling formula with Perturb-and-Average and variational method [93, 89], improved noise sampling schedules [34], various 3D representations [46, 87, 11], text prompts [2], 3D consistency [44, 32], and prior quality with dedicated diffusion models [69, 78]. Our method leverages the Normal-Depth diffusion model used in [69] for text-to-3D shape generation, which is trained on the large-scale LAION dataset, but replaces the geometry representation with our TetSphere.

### v4

@@ Appendix I More Implementation Details
The recent success of text-to-2D image generation models has spurred a growing interest in generating 3D output from text input. In light of the limited availablity of text-annotated 3D datasets, methods have been developed that leverage the pre-trained text-to-image models to reason between 2D renderings of 3D models and text descriptions. Early works (Khalid et al., 2022; Jain et al., 2022) adopt the pre-trained CLIP model (Radford et al., 2021) to supervise the generation by aligning the clip text and image embeddings. Recently, 2D diffusion-based generative models (Saharia et al., 2022; Rombach et al., 2022) have powered direct supervision in the image/latent space and achieved superior 3D quality (Poole et al., 2022; Tang et al., 2023b; Chen et al., 2023; Lin et al., 2023; Wang et al., 2023c). Notably, DreamFusion (Poole et al., 2022) introduces the Score Distillation Sampling (SDS) for supervising the NeRF (Mildenhall et al., 2020) optimization using diffusion priors as a score function (i.e., by minimizing the added noise and the predicted noise under the text condition). Follow-up works have since been proposed to improve score sampling formula with Perturb-and-Average and variational method (Wang et al., 2023c; a), improved noise sampling schedules (Huang et al., 2023), various 3D representations (Lin et al., 2023; Tsalicoglou et al., 2023; Chen et al., 2023), text prompts (Armandpour et al., 2023), 3D consistency (Li et al., 2023; Hong et al., 2023a), and prior quality with dedicated diffusion models (Qiu et al., 2023; Shi et al., 2023). Our method leverages the Normal-Depth diffusion model used in (Qiu et al., 2023) for text-to-3D shape generation, which is trained on the large-scale LAION dataset, but replaces the geometry representation with our TetSphere.
Appendix K Limitations
While our representation demonstrates improvements in reconstruction quality, there are notable limitations that should be acknowledged. For example, the method only produces piecewise meshes that are overlaid, rather than solving the problem of globally fitting a single tet-mesh. Furthermore, while our regularization helps prevent issues such as tetrahedron inversion and maintains mesh integrity, it does not provide strong topological guarantees. The union operation of multiple TetSpheres can potentially alter the topology and may not fully preserve manifold properties. Addressing these challenges to ensure global mesh consistency and topological guarantees could be valuable directions for future work.
Appendix L Discussions on Bi-harmonic Energy on Deformation Gradients
The incorporation of biharmonic energy in our method is inspired by its proven efficacy in geometry processing, as highlighted by prior research (Botsch & Sorkine, 2007). Furthermore, penalizing the non-smoothness of transformations (or deformation gradient) across a domain is a well-explored area that has demonstrated its utility in various geometry processing tasks such as shape deformation (Wang et al., 2021a), deformation transfer (Sumner & Popović, 2004), and mesh alignment Amberg et al. (2007). For the surface regions where observations are available, the surface deformation is primarily influenced by the rendering loss. Conversely, in regions lacking direct observations, our regularization term plays an important role by penalizing high-frequency deformations resulting from non-smooth deformation gradients. This approach ensures reliable results. The biharmonic energy offers stronger regularization compared to first-order harmonic energy, as demonstrated in (Botsch & Sorkine, 2007). Although theoretically higher-order smoothness energies can be considered, in practice, they often lead to numerical instability, as discussed in Jacobson & Panozzo (2017). Therefore, biharmonic energy strikes a balance by providing effective regularization while avoiding numerical issues.